# The long way of a viscous vortex dipole

### v1

@@ 4.8. The forcing term
Let $\mathrm{F}$ be the forcing term (4.55). Under the assumptions of Proposition 4.10, for any $\kappa_{*}>0$, there exists a constant $C_{*}>0$ such that
@@ 4.9. The nonlinear term
Let $\mathrm{NL}$ be the nonlinearity defined in (4.56). Under the assumptions of Proposition 4.10, there exist constants $C_{*}>0$ and $10<s_{*}<$ 1 such that
@@
To estimate the right-hand side, we consider separately the regions $\mathrm{I}_{\varepsilon}$ and $\mathrm{I}_{\varepsilon}^{c}$. Applying Hölder’s inequality with $21/q+1/p=1/$ 2 and $q>2$, and using Lemmas 4.54.6 to bound the stream function $\mathcal{T}_{\varepsilon}\varphi$, we obtain
$\|\mathbbm{1}_{\rm I_{\varepsilon}}W_{\varepsilon}^{-\frac{1}{2}}\nabla\mathcal{T}_{\varepsilon}\varphi\|_{L^{2}}\,$ $\lesssim\,\|\mathbbm{1}_{\rm I_{\varepsilon}}(1+|\cdot|)^{3}W_{\varepsilon}^{-\frac{1}{2}}\|_{L^{p}}\|\mathbbm{1}_{\rm I_{\varepsilon}}(1+|\cdot|)^{-3}\nabla\mathcal{T}_{\varepsilon}\varphi\|_{L^{q}}\,\lesssim\,\varepsilon^{3}\|w\|_{\mathcal{X}_{\varepsilon}}\,,$
@@
Here we use the bound (4.32) in Lemma 4.5 to control the streamfunction, and we make the following observations concerning $\nabla W_{\varepsilon}$. In region $\mathrm{I}_{\varepsilon}$, thanks to Proposition 4.4, we can approximate $\nabla W_{\varepsilon}$ by $\nabla W_{0}$, which gives $|\nabla W_{\varepsilon}(\xi)|\lesssim|\nabla W_{0}(\xi)|+W_{0}(\xi)\lesssim\rho_{\varepsilon}(\xi)W_{\varepsilon}(\xi)$, for $|\xi|\leq 2\varepsilon^{-\sigma_{1}}$. In region $\mathrm{II}_{\varepsilon}$ we have $\nabla W_{\varepsilon}=0$ whereas in region $\mathrm{III}_{\varepsilon}$ we know that $|\nabla W_{\varepsilon}|\lesssim|\xi|^{\gamma-1}\rho_{\varepsilon}W_{\varepsilon}$. Overall, we obtain
|$\begin{split}&\bigl{|}\left\langle\left\{\varphi-\mathcal{T}_{\varepsilon}\varphi,W_{\varepsilon}\right\}w,w\right\rangle_{L^{2}}\bigr{|}\,\lesssim\,\|\nabla\varphi\|_{L^{\infty}}\|\rho_{\varepsilon}w\|_{\mathcal{X}_{\varepsilon}}\|w\|_{\mathcal{X}_{\varepsilon}}\\ &\qquad\,\lesssim\,\bigl{(}\|\nabla w\|_{\mathcal{X}_{\varepsilon}}^{\frac{1}{2}}+\|w\|^{\frac{1}{2}}_{\mathcal{X}_{\varepsilon}}\bigr{)}\|\rho_{\varepsilon}w\|_{\mathcal{X}_{\varepsilon}}\|w\|^{\frac{3}{2}}_{\mathcal{X}_{\varepsilon}}\,\lesssim\,\sqrt{E_{\varepsilon}[w]}D_{\varepsilon}[w]\,,\end{split}$ (4.116)
@@
Under the assumptions of Proposition 4.10, we have $\delta^{-1}\varepsilon^{M+1}+\delta\varepsilon^{2}\leq\delta^{s_{*}}$ for some $s_{*}>0$, see Remark 4.11. Therefore, combining the the bounds (4.115), (4.116) and (4.117), we see that the nonlinearity $\mathrm{NL}$ defined by (4.56) satisfies estimate (4.114) for some constant $C_{*}>0$. ∎
@@ A.2. Inverting the diffusion operator
$(\kappa-\mathcal{L})^{-1}f\,=\,\int_{0}^{\infty}e^{-\kappa\tau}S(\tau)f\,{\rm d}\tau\,,$ (A.4)
where $S(\tau)=\exp(\tau\mathcal{L})$ is the analytic semigroup in $\mathcal{Y}$ generated by the selfadjoint operator $\mathcal{L}$, see [15, Appendix A]. It is well known that
@@
$\frac{1}{4}\bigl{(}|\xi|^{2}-|\eta|^{2}\bigr{)}-\frac{1}{4a(\tau)}\,|\xi-\eta\,{\rm e}^{-\tau/2}|^{2}\,=\,-\frac{1}{4a(\tau)}\,|\eta-\xi\,{\rm e}^{-\tau/2}|^{2}\,,$
@@
${\rm e}^{|\xi|^{2}/4}\bigl{|}\bigl{(}(\kappa-\mathcal{L})^{-1}f\bigr{)}(\xi)\bigr{|}\,\leq\,\int_{0}^{\infty}e^{-\kappa\tau}|G(\xi,\tau)|\,{\rm d}\tau\,\leq\,C\kappa^{-1}(1+|\xi|)^{N}\,.$ (A.6)
The derivative $\partial^{\alpha}(\kappa-\mathcal{L})^{-1}f$ can be estimated in the same way, for any multi-index $\alpha\in\mathbb{N}^{2}$. If $10<\tau<$ 1 we differentiate (A.5) and integrate by parts to obtain
@@
Since $f\in\mathcal{Z}$ we have $e^{|\eta|^{2}/4}|\partial_{\eta}^{\alpha}f(\eta)|\leq C(1+|\eta|)^{N^{\prime}}$ for some $N^{\prime}\in\mathbb{N}$ (depending on $\alpha$), and proceeding as before we deduce that $e^{|\xi|^{2}/4}|\partial_{\xi}^{\alpha}\bigl{(}S(\tau)f\bigr{)}(\xi)|\leq C(1+|\xi|)^{N^{\prime}}$. If $\tau\geq 1$, we observe that $1-{\rm e}^{-1}\leq a(\tau)\leq 1$ and we differentiate (A.5) to obtain
@@
which gives $e^{|\xi|^{2}/4}|\partial_{\xi}^{\alpha}\bigl{(}S(\tau)f\bigr{)}(\xi)|\leq C(1+|\xi|)^{N+|\alpha|}$. Thus using (A.4) we find, as in (A.6),
${\rm e}^{|\xi|^{2}/4}\bigl{|}\partial^{\alpha}\bigl{(}(\kappa-\mathcal{L})^{-1}f\bigr{)}(\xi)\bigr{|}\,\leq\,\int_{0}^{\infty}e^{-\kappa\tau}{\rm e}^{|\xi|^{2}/4}\bigl{|}\partial_{\xi}^{\alpha}\bigl{(}S(\tau)f\bigr{)}(\xi)\bigr{|}\,{\rm d}\tau\,\leq\,C\kappa^{-1}(1+|\xi|)^{N^{\prime\prime}}\,,$ (A.7)
@@ A.3. Inverting the advection operator
$-\varphi^{\prime\prime}(r)-\frac{1}{r}\,\varphi^{\prime}(r)+\Bigl{(}\frac{n^{2}}{r^{2}}-h(r)\Bigr{)}\varphi(r)\,=\,\frac{b(r)}{nv_{0}(r)}\,,\qquad r>0\,,$ (A.11)
@@
A.4. Proof of Lemma 4.6
Since $\mathrm{M}[w]=0$, using (2.6) and the definition of $\mathcal{T}_{\varepsilon}$ in (2.10) we find as in Lemma 3.8
$\bigl{(}\mathcal{T}_{\varepsilon}\varphi\bigr{)}(\xi)\,$ $=\,\frac{1}{4\pi}\int_{\mathbb{R}^{2}}\log\left(|1+\varepsilon(\xi_{1}+\eta_{1})|^{2}+\varepsilon^{2}|\xi_{2}-\eta_{2}|^{2}\right)w(\eta)\,{\rm d}\eta$
$=:\,\int_{\mathbb{R}^{2}}K_{\varepsilon}(\xi_{1}+\eta_{1},\xi_{2}-\eta_{2})w(\eta)\,{\rm d}\eta\,,\qquad\xi\in\mathbb{R}^{2}\,.$
We first prove the bound on $\nabla\mathcal{T}_{\varepsilon}\varphi$ in (4.34). To evaluate the contribution of the vorticity $w$ in the outer region, we define
$I_{\mathrm{out}}(\xi)\,:=\,\int_{\{|\eta|>4\varepsilon^{-\sigma_{1}}\}}\nabla_{\xi}K_{\varepsilon}(\xi_{1}+\eta_{1},\xi_{2}-\eta_{2})w(\eta)\,{\rm d}\eta\,,\qquad\xi\in\mathbb{R}^{2}\,.$
Given any $q>2$, we claim that
$\|I_{\mathrm{out}}\|_{L^{q}}\,\lesssim\,\exp(-c_{*}\varepsilon^{-2\sigma_{1}})\|w\|_{\mathcal{X}_{\varepsilon}}\,,$ (A.14)
for some sufficiently small constant $c_{*}>0$. To this end, we observe that
$|\nabla_{\xi}K_{\varepsilon}(\xi_{1}+\eta_{1},\xi_{2}-\eta_{2})|\,\lesssim\,\frac{1}{|\varepsilon^{-1}+(\xi_{1}+\eta_{1})|+|\xi_{2}-\eta_{2}|}\,.$
Using the change of variables $\eta_{1}=-\varsigma_{1}-\varepsilon^{-1},\,\eta_{2}=\varsigma_{2}$, we thus get
$\int_{\{|\eta|>4\varepsilon^{-\sigma_{1}}\}}|\nabla_{\xi}K_{\varepsilon}(\xi_{1}+\eta_{1},\xi_{2}-\eta_{2})w(\eta)|\,{\rm d}\eta\,\lesssim\,\int_{\mathbb{R}^{2}}\frac{1}{|\xi-\varsigma|}\,\mathcal{T}_{\varepsilon}\bigl{(}\mathbbm{1}_{\{|\cdot|>4\varepsilon^{-\sigma_{1}}\}}|w|\bigr{)}(\varsigma){\rm d}\varsigma\,.$
Taking $p\in(1,2)$ such that $21/p=1/q+1/$ 2 and applying the Hardy–Littlewood–Sobolev inequality, we deduce that
$\|I_{\mathrm{out}}\|_{L^{q}}\,\lesssim\,\biggl{(}\int_{\{|\eta|>4\varepsilon^{-\sigma_{1}}\}}W_{\varepsilon}^{-\frac{p}{2}}(\eta)\bigl{(}\sqrt{W_{\varepsilon}}|w|\bigr{)}^{p}(\eta)\,{\rm d}\eta\biggr{)}^{\frac{1}{p}}\,\lesssim\,\|\mathbbm{1}_{\{|\cdot|>4\varepsilon^{-\sigma_{1}}\}}W_{\varepsilon}^{-\frac{1}{2}}\|_{L^{\frac{2p}{2-p}}}\|w\|_{\mathcal{X}_{\varepsilon}}\,,$
where the last step follows from Hölder’s inequality with conjugate exponents $2/(2{-}p)$ and $2/p$. Finally, using the lower bound on $W_{\varepsilon}$ given by the definition (4.22), we arrive at the bound (A.14) for some sufficiently small $c_{*}>0$.
We next consider the contribution of the vorticity $w(\eta)$ in the complementary region where $|\eta|\leq 4\varepsilon^{-\sigma_{1}}$. We assume henceforth that $\xi\in\mathrm{I}_{\varepsilon}$, so that $|\xi|\leq 2\varepsilon^{-\sigma_{1}}$. This of course implies that $\varepsilon(|\xi|+|\eta|)\ll 1$, so that we can expand the kernel $K_{\varepsilon}(\xi_{1}+\eta_{1},\xi_{2}-\eta_{2})$ into a convergent power series as in (3.19). In particular, if we define
$B_{\mathrm{err}}(\xi,\eta)\,:=\,\nabla_{\xi}K_{\varepsilon}(\xi_{1}+\eta_{1},\xi_{2}-\eta_{2})-\sum_{n=1}^{2}\frac{(-1)^{n-1}}{2\pi n}\,\varepsilon^{n}\nabla_{\xi}Q_{n}^{c}(\xi_{1}+\eta_{1},\xi_{2}-\eta_{2})\,,$
we have the accurate bound $|B_{\mathrm{err}}(\xi,\eta)|\lesssim\varepsilon^{3}(1+|\xi|+|\eta|)^{2}$. Moreover, by construction,
$\begin{split}\bigl{(}\nabla\mathcal{T}_{\varepsilon}\varphi\bigr{)}(\xi)\,=\,&\sum_{n=1}^{2}\frac{(-1)^{n-1}}{2\pi n}\,\varepsilon^{n}\nabla_{\xi}\int_{\{|\eta|\leq 4\varepsilon^{-\sigma_{1}}\}}Q_{n}^{c}(\xi_{1}+\eta_{1},\xi_{2}-\eta_{2})w(\eta)\,{\rm d}\eta\\ &+\int_{\{|\eta|\leq 4\varepsilon^{-\sigma_{1}}\}}B_{\mathrm{err}}(\xi,\eta)w(\eta)\,{\rm d}\eta+I_{\mathrm{out}}(\xi)\,,\quad\text{for }\,|\xi|\leq 2\varepsilon^{-\sigma_{1}}\,.\end{split}$ (A.15)
In view of the definition of the polynomials $P_{n}(\xi)$ in (3.17) (with $\Omega=w$), the first term in the right-hand side of (A.15) is equal to
$\nabla P_{1}(\xi)+\nabla P_{2}(\xi)-\sum_{n=1}^{2}\frac{(-1)^{n-1}}{2\pi n}\varepsilon^{n}\nabla_{\xi}\int_{\{|\eta|>4\varepsilon^{-\sigma_{1}}\}}Q_{n}^{c}(\xi_{1}+\eta_{1},\xi_{2}-\eta_{2})w(\eta)\,{\rm d}\eta\,.$
As observed in Remark 3.9, we know that $\nabla P_{1}(\xi)=\nabla P_{2}(\xi)=0$ if $\mathrm{M}[w]=\mathrm{m}_{1}[w]=\mathrm{m}_{0}[w]=0$. Therefore, since $|\nabla Q_{n}^{c}(x)|\leq n|x|^{n-1}$, we get

### v2

@@ 4.8. The forcing terms
Let $\mathrm{F}$ be defined as in (4.56). Under the assumptions of Proposition 4.11, for any $\kappa_{*}>0$, there exists a constant $C_{*}>0$ such that
@@ 4.9. The nonlinear terms
Let $\mathrm{NL}$ be defined as in (4.57). Under the assumptions of Proposition 4.11, there exist constants $C_{*}>0$ and $0<s_{*}<1$ such that
@@
To estimate the right-hand side, we consider separately the regions $\mathrm{I}_{\varepsilon}$ and $\mathrm{I}_{\varepsilon}^{c}$. Applying Hölder’s inequality with $1/q+1/p=1/2$ and $q>2$, and using Lemma 4.5 and Corollary 4.7 to bound the stream function $\mathcal{T}_{\varepsilon}\varphi$, we obtain
$\|\mathbbm{1}_{\rm I_{\varepsilon}}W_{\varepsilon}^{-\frac{1}{2}}\nabla\mathcal{T}_{\varepsilon}\varphi\|_{L^{2}}\,$ $\lesssim\,\|\mathbbm{1}_{\rm I_{\varepsilon}}(1+|\cdot|)W_{\varepsilon}^{-\frac{1}{2}}\|_{L^{p}}\|\mathbbm{1}_{\rm I_{\varepsilon}}(1+|\cdot|)^{-1}\nabla\mathcal{T}_{\varepsilon}\varphi\|_{L^{q}}\,\lesssim\,\varepsilon^{3}\|w\|_{\mathcal{X}_{\varepsilon}}\,$,
@@
Here we use the bound (4.32) in Lemma 4.5 to control the streamfunction, and we make the following observations concerning $\nabla W_{\varepsilon}$. In region $\mathrm{I}_{\varepsilon}$, thanks to Proposition 4.4, we can approximate $\nabla W_{\varepsilon}$ by $\nabla W_{0}$, which gives $|\nabla W_{\varepsilon}(\xi)|\lesssim|\nabla W_{0}(\xi)|+W_{0}(\xi)\lesssim(1+\rho_{\varepsilon}(\xi))W_{\varepsilon}(\xi)$, for $|\xi|\leq 2\varepsilon^{-\sigma_{1}}$. In region $\mathrm{II}_{\varepsilon}$ we have $\nabla W_{\varepsilon}=0$ whereas in region $\mathrm{III}_{\varepsilon}$ we know that $|\nabla W_{\varepsilon}|\lesssim|\xi|^{\gamma-1}\rho_{\varepsilon}W_{\varepsilon}$. Overall, we obtain
[…]
@@
Under the assumptions of Proposition 4.11, we have $\delta^{-1}\varepsilon^{M+1}+\delta\varepsilon^{2}\leq\delta^{s_{*}}$ for some $s_{*}>0$, see Remark 4.12. Therefore, combining the the bounds (4.115), (4.116) and (4.117), we see that $\mathrm{NL}$ defined by (4.57) satisfies the estimate (4.114) for some constant $C_{*}>0$. ∎
@@ A.2. Inverting the diffusion operator
$(\kappa-\mathcal{L})^{-1}f\,=\,\int_{0}^{\infty}{\rm e}^{-\kappa\tau}S(\tau)f\,{\rm d}\tau\,$, (A.4)
where $S(\tau)=\exp(\tau\mathcal{L})$ is the analytic semigroup in $\mathcal{Y}$ generated by the selfadjoint operator $\mathcal{L}$, see [16, Appendix A]. It is well known that
@@
$\frac{|\xi|^{2}}{4}-\frac{1}{4a(\tau)}\,\bigl|\xi-\eta\,{\rm e}^{-\tau/2}\bigr|^{2}\,=\,\frac{|\eta|^{2}}{4}-\frac{1}{4a(\tau)}\,\bigl|\eta-\xi\,{\rm e}^{-\tau/2}\bigr|^{2}\,$,
@@
${\rm e}^{|\xi|^{2}/4}\bigl|\bigl((\kappa-\mathcal{L})^{-1}f\bigr)(\xi)\bigr|\,\leq\,\int_{0}^{\infty}{\rm e}^{-\kappa\tau}|G(\xi,\tau)|\,{\rm d}\tau\,\leq\,C\kappa^{-1}(1+|\xi|)^{N}\,.$ (A.6)
The derivative $\partial^{\alpha}(\kappa-\mathcal{L})^{-1}f$ can be estimated in the same way, for any multi-index $\alpha\in\mathbb{N}^{2}$. If $0<\tau<1$ we differentiate (A.5) and integrate by parts to obtain
@@
Since $f\in\mathcal{Z}$ we have ${\rm e}^{|\eta|^{2}/4}|\partial_{\eta}^{\alpha}f(\eta)|\leq C(1+|\eta|)^{N^{\prime}}$ for some $N^{\prime}\in\mathbb{N}$ (depending on $\alpha$), and proceeding as before we deduce that ${\rm e}^{|\xi|^{2}/4}|\partial_{\xi}^{\alpha}\bigl(S(\tau)f\bigr)(\xi)|\leq C(1+|\xi|)^{N^{\prime}}$. If $\tau\geq 1$, we observe that $1-{\rm e}^{-1}\leq a(\tau)\leq 1$ and we differentiate (A.5) to obtain
@@
which gives ${\rm e}^{|\xi|^{2}/4}|\partial_{\xi}^{\alpha}\bigl(S(\tau)f\bigr)(\xi)|\leq C(1+|\xi|)^{N+|\alpha|}$. Thus using (A.4) we find, as in (A.6),
${\rm e}^{|\xi|^{2}/4}\bigl|\partial^{\alpha}\bigl((\kappa-\mathcal{L})^{-1}f\bigr)(\xi)\bigr|\,\leq\,\int_{0}^{\infty}{\rm e}^{-\kappa\tau}{\rm e}^{|\xi|^{2}/4}\bigl|\partial_{\xi}^{\alpha}\bigl(S(\tau)f\bigr)(\xi)\bigr|\,{\rm d}\tau\,\leq\,C\kappa^{-1}(1+|\xi|)^{N^{\prime\prime}}\,$, (A.7)
@@ A.3. Inverting the advection operator
$\varphi^{\prime\prime}(r)-\frac{1}{r}\,\varphi^{\prime}(r)+\Bigl(\frac{n^{2}}{r^{2}}-h(r)\Bigr)\varphi(r)\,=\,\frac{b(r)}{nv_{0}(r)}\,,\qquad r>0\,$, (A.11)
@@
A.4. Estimates for the stream function
In this final section we give a short proof of Lemmas 4.54.6 and of Corollary 4.7.
Proof of Lemma 4.5.
In view of (2.6) we have, for all $\xi\in\mathbb{R}^{2}$,